A box rests on top of a flat bed truck. The box ha
Answer #1

1. Vf=vi + a(t)
a=19/12=1.58

2. F=Ma=(20*1.58)=31.6

3. Normal force= mg = (9.8)(20)=196
Force=UsN
Ff=(.86)(196)=168.56
168.56=Ma
A=168.56/20=8.428

4. Ff=UkN
Ff=(.67)(196)
Ff=131.32
131.32/20=6.566

5. Same answer as Number 3